OpenStudy (anonymous):

show that if |z-5i/z+5i|=1 therfore z is a real number

OpenStudy (experimentx):

let z = x+iy, \[ |x + i y - 5i| = |x + i y + 5i| \\ x^2 + (y-5)^2 = x^2 + (y+5)^2 \implies y = 0\]

OpenStudy (anonymous):

is this a compltete answer

OpenStudy (experimentx):

you need to show that the imaginary part is zero.

OpenStudy (anonymous):

see..\[\left| z-5i/z+5i \right|=\left| z-5i \right|/\left| z+5i \right|=1\] now..\[\left| z-5i \right|=\left| z+5i \right| \] which means that z is a complex no. which is equi distant from both 5i and -5i...SO Z IS AREAL NUMBER